# The local converse theorem for $\widetilde{\mathrm{Sp}}_{2n}$

### v5

@@ 1. Introduction
Let $F$ be a non-archimedean local field of characteristic not equal to 2 (i.e., a finite extension of $\mathbb{Q}_{p}$ for any prime $p$ or $\mathbb{F}_{q}((t))$ for odd prime power $q=p^{k}$), and let $W$ be a $2n$-dimensional symplectic space over $F$. Define $\widetilde{\mathrm{J}_{n}}=\widetilde{\mathrm{Sp}}(W)$ as the metaplectic group of $W$, which is the double covering group of $\mathrm{J}_{n}=\mathrm{Sp}(W)$, the isometry group of $W$.
We introduce $\psi$, a non-trivial additive character of $F$. For $\lambda\in F^{\times}$, we define the character $\psi_{\lambda}$ of $F$ as $\psi_{\lambda}(x)\coloneqq\psi(\lambda x)$. Let $\mathrm{U}^{\prime}$ be the unipotent radical of a Borel subgroup of $\mathrm{Sp}(W)$, and let $\mu_{\lambda}^{\prime}$ be a certain non-trivial character of $\mathrm{U}^{\prime}(F)$ that depends on $\psi$ and $\lambda$ (see Section 2.2 for its definition).
@@
Let $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ be irreducible admissible $\mu_{\lambda}^{\prime}$-generic representations of $\widetilde{\mathrm{J}_{n}}(F)$, satisfying the condition
@@
Around 2017, Chai [Cha19] and Jacquet and Liu [JL18] independently refined Henniart’s result to its optimal form.
More recently, Q. Zhang [Q18, Q19] proved the supercuspidal cases of the conjecture for $\mathrm{Sp}_{2n}$ and $\mathrm{U}_{2n+1}$ using the theory of partial Bessel functions developed by Cogdell, Shahidi, and Tsai in [CST17]. In a similar vein, B. Liu and A. Hazeltine [LH22] established the conjecture for split even special orthogonal groups, while Y. Jo [Jo] refined Jiang and Soudry’s [JS03] result and extended Zhang’s [Q18] result to the generic cases. Notably, Jo’s work addressed local fields of positive characteristics as well as characteristic zero, in contrast to other LCT results that only dealt with characteristic zero cases. These proofs were purely local and did not invoke the weak Langlands functorial lifts to general linear groups, as in the recent results of [Cha19] and [JL18]. For other classical groups, Morimoto [M] demonstrated the theorem for even unitary groups by pulling back the recent results of [Cha19] and [JL18] using global (and hence) local descent methods.
@@
To successfully transfer the LCT, it is necessary to demonstrate that the LTC preserves certain properties and invariants of representations. Specifically, we consider genericity (Theorem 3.3) and $\gamma$-factors (Theorem 3.7). Indeed, these results can be found in [GS, Theorem 9.1] and [GS, Proposition 11.1] accompanied by [GS, Corollary 11.2], respectively.
However, unfortunately, the proof of [GS, Theorem 9.1] is absent in the cited reference. To address this gap, we provide a complete proof demonstrating that the missing argument holds even in the case when $\operatorname{char}(F)=p$. Additionally, [GS, Proposition 11.1] relies on the general $\gamma$-factor theory, which is currently unavailable. Consequently, we furnish a comprehensive proof for this proposition using the established generic $\gamma$-factor theory (see Remark 3.8).
Theorem 3.3 and Theorem 3.7 form the main part of the paper and are discussed in Section 3, along with Section 4, which computes the twisted Jacquet module of the Weil representation—a crucial step in proving the preservation of genericity.”
@@
When the residual characteristic of the $p$-adic field $F$ is odd, Zhang ([Q17]) directly proved this result under the condition that $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ share the same central characters. Our result covers cases where the residual characteristic of $F$ is 2, and it also extends to the situation when $F=\mathbb{F}_{q}((t))$, without requiring the assumption of identical central characters. Alternatively, we can remove the assumption on the same central characters from the local functional equation of doubling zeta integrals of metaplectic groups (see [GS, Section 4.3]).
@@ 2.3. $\gamma$-factors
Let $\mathrm{G}$ be either $\mathrm{H}_{m}=\mathrm{SO}_{2m+1}$ or $\widetilde{\mathrm{J}_{n}}=\widetilde{\mathrm{Sp}}_{2n}$ Let $\pi$ be an irreducible generic representation of $\mathrm{G}(F)$ and $\sigma$ an irreducible generic representation of $\mathrm{GL}_{l}(F)$. The local $\gamma$-factors can be defined in two distinct ways. The first approach is through the Langlands–Shahidi method ([Sha90], [L15], [Sz10]), while the second approach relies on the proportionality between the functional equations of local Rankin–Selberg type integrals (or Shimura type integrals) for $\mathrm{H}_{m}\times\mathrm{GL}_{l}$ (or $\widetilde{\mathrm{J}_{n}}\times\mathrm{GL}_{l}$). For a precise definition, please refer to [So93] and [Kap15]. It is worth noting that the two definitions of $\gamma$-factors are reconciled by several properties, as explained in [Kap15]. Since we consider both the cases of $\operatorname{char}(F)=0$ and $\operatorname{char}(F)=p$ simultaneously, we adopt the definition of $\gamma$-factors from the Langlands–Shahidi method, as they are defined in both cases.
We would like to recall some important properties of the $\gamma$-factors that will be utilized later in our discussion.
Proposition 2.1 ([Kap15], Theorem 1, [Sha90], Theorem 3.5, [L15], Theorem 1.4).
@@ 7. $(\mathrm{vii})$
    (Tempered $L$-functions) Suppose that $\pi,\sigma$ are irreducible tempered representations of $G(F)$ and $\mathrm{GL}_{l}(F)$. Let $P_{\pi\times\sigma}(t)$ be the unique polynomial with $P_{\pi\times\sigma}(0)=1$ such that $P_{\pi\times\sigma}(q^{-s})$ is the numerator of $\gamma(s,\pi\times\sigma,\psi)$. Then $L_{\psi}(s,\pi\times\sigma)$ is defined by
    $L_{\psi}(s,\pi\times\sigma)\coloneqq P_{\pi\times\sigma}(q^{-s})^{-1}.$
  8. $(\mathrm{viii})$
    (Holomorphy of tempered $\epsilon$-factors) Suppose that $\pi,\sigma$ are irreducible tempered representations of $G(F)$ and $\mathrm{GL}_{l}(F)$. Then
    $\epsilon(s,\pi\times\sigma,\psi)\coloneqq\gamma(s,\pi\times\sigma,\psi)\frac{L_{\psi}(s,\pi\times\sigma,\psi)}{L_{\psi}(1-s,\pi^{\vee}\times\sigma^{\vee},\psi)}$
    is a monomial of $q^{-s}$.
@@ 8. $(\mathrm{viii})$
The local unramified $L$-function $L_{\psi}(s,\pi\times\sigma)$ appearing in (iii) is defined as follows. Let $\pi$ be the principal series representation corresponding to the (genuine) unramified character
$\mathrm{diag}(t_{1},\cdots,t_{m},t_{m}^{-1},\cdots,t_{1})\mapsto\prod_{i=1}^{m}\mu_{i}(t_{i})\quad\text{ or }\quad\left(\mathrm{diag}(t_{1},\cdots,t_{n},t_{n}^{-1},\cdots,t_{1}),\epsilon\right)\mapsto\epsilon\cdot\gamma_{\psi}\left(\prod_{i=1}^{n}t_{i}\right)\prod_{i=1}^{n}\mu_{i}(t_{i}).$
Also, assume that $\sigma$ is the principal series representation induced from the unramified character $\bigotimes_{i=1}^{l}\chi_{i}$ of the maximal torus of $\mathrm{GL}_{l}(F)$. Then, $L_{\psi}(s,\pi\times\sigma)$ is defined as
$\prod_{i,j}\left(\frac{1}{1-\mu_{i}(\varpi)\chi_{j}(\varpi)q^{-s}}\right)\cdot\left(\frac{1}{1-\mu_{i}(\varpi)^{-1}\chi_{j}(\varpi)q^{-s}}\right).$
Note that $\psi$ in $L_{\psi}$ appears only when dealing with metaplectic groups.
Remark 2.3.
While the definition and standard properties of $\gamma$-factors for metaplectic groups in characteristic $\operatorname{char}(F)=p$ have not yet been firmly established, it is reasonable to expect that the methods introduced in [L15] naturally extend to this case. However, due to the focus and limitations of this paper, we will adopt the working hypothesis that such extensions exist without delving into the details.
@@
The $\gamma$-factors for $\widetilde{\mathrm{J}_{n}}\times\mathrm{GL}_{l}$ are properly defined in $\operatorname{char}(F)=p$ cases. Furthermore, they satisfy properties (i)-(viii) in Proposition 2.1.
@@ 3. Local theta correspondence
For the rest of this work, we focus exclusively on the case where $m=n$ and conveniently denote $V_{n}$ and $W_{n}$ as $V$ and $W$ respectively. Given $\widetilde{\pi}\in\operatorname{Irr}(\widetilde{\mathrm{J}_{n}})$, there exists a unique $\epsilon\in\{\pm 1\}$ such that $\theta_{\psi,V^{\epsilon},W}(\widetilde{\pi})$ is nonzero (see [GS, Theorem 1.4]).
@@ 3. $(\mathrm{iii})$
When $r=1$, the theorem is incorporated into [GI1, Theorem 11.5] in $\operatorname{char}(F)=0$ case. In the case of $\operatorname{char}(F)=p\neq 2$, a parallel reasoning to the proof of [GI1, Theorem 11.5] can be applied, utilizing the global-to-local argument from [GI1, subsection 11.6]. Notably, in this scenario, [GL, Theorem 1.1] assumes a comparable role. Consequently, we may presume the validity of [GI1, Theorem 11.5] across all characteristics except 2.
@@
Let $\Xi$ be the global theta lift of $\Pi$ to $\mathrm{SO}(\mathbb{V})(\mathbb{A})$. Using [Fu, Proposition 3], we know that $\Xi$ is nonzero and generic. We claim that $\Xi$ is cuspidal. Suppose not, which means that there exists an integer $m<n$ and a $(2m+1)$-dimensional symmetric space $\mathbb{V}_{m}$ over $K$ such that the global theta lift $\Theta_{m}(\Pi)$ of $\Pi$ to $\mathrm{SO}(\mathbb{V}_{m})(\mathbb{A})$ is cuspidal. Let $\Theta_{m}(\Pi)_{0}$ be an irreducible summand of $\Theta_{m}(\Pi)$. Then $\Theta_{m}(\Pi)_{0,v_{0}}$ is the nonzero local theta lift of $\widetilde{\pi}$ to $\mathrm{SO}(V_{m})$. However, this contradicts the tower property of local theta lifts, since $\pi$ is a supercuspidal representation of $\mathrm{H}_{n}(F)$. Therefore, we have shown that $\Xi$ is a cuspidal representation of $\mathrm{SO}(\mathbb{V})(\mathbb{A})$ with $\Xi_{v_{0}}=\theta_{\psi,V,W}(\widetilde{\pi})$.
@@ 5. The proof
For an irreducible generic tempered representation $\pi$ of $\mathrm{H}_{n}(F)$ and unitary supercuspidal representation $\sigma$ of $\mathrm{GL}_{k}(F)$, the followings hold.
@@ 1. $(\mathrm{i})$
    $\gamma(s,\pi\times\sigma,\psi)$ has no zero for $\operatorname{Re}(s)>0$.
@@ 2. $(\mathrm{ii})$
    If $\gamma(s,\pi\times\sigma,\psi)$ has a real pole at $s=s_{0}$, then the pole must be a simple pole at $s_{0}=1$ and $\sigma\simeq\sigma^{\vee}$.
@@
By Proposition 2.1 (vii), we have
$\gamma(s,\pi\times\sigma,\psi)=\epsilon(s,\pi\times\sigma,\psi)\cdot\frac{L(1-s,\pi^{\vee}\times\sigma^{\vee},\psi)}{L(s,\pi\times\sigma,\psi)}.$
According to [CS, Corollary 4.2] and [L19, Corollary 5.5], $L(s,\pi\times\sigma,\psi)$ is holomorphic for $\operatorname{Re}(s)>0$ and so property $(\mathrm{i})$ follows from this observation and the holomorphy of tempered $\epsilon$-factors (Proposition 2.1 (vii).)
In the proof of [Jo, Proposition 4.12 $(\mathrm{iii})$], the requirement that $\pi$ is supercuspidal was imposed in order to apply the ’Casselman–Shahidi’ Lemma ([Jo, Lemma 4.10]). However, it has been established in [Kim, Proposition 12.3] that the ’Casselman–Shahidi Lemma’ also holds for tempered representations. Therefore, except for this specific assumption, all other arguments presented in [Jo, Proposition 4.12 $(\mathrm{iii})$] can be equally applied to tempered representations. Hence, property $(\mathrm{ii})$ is proven. ∎

### v6

@@ 1. Introduction
Let $F$ be a non-archimedean local field of characteristic not equal to 2 (i.e., a finite extension of $\mathbb{Q}_{p}$ for any odd prime $p$ or $\mathbb{F}_{q}((t))$ for $q=p^{k}$), and let $W$ be a $2n$-dimensional symplectic space over $F$. Define $\widetilde{\mathrm{J}_{n}}=\widetilde{\mathrm{Sp}}(W)$ as the metaplectic group of $W$, which is the double covering group of $\mathrm{J}_{n}=\mathrm{Sp}(W)$, the isometry group of $W$.
We introduce $\psi$, a non-trivial additive character of $F$. For $\lambda\in F^{\times}$, we define the character $\psi_{\lambda}$ of $F$ as $\psi_{\lambda}(x)\coloneqq\psi(\lambda x)$. Let $\mathrm{U}^{\prime}$ be the unipotent radical of a Borel subgroup of $\mathrm{Sp}(W)$, and let $\mu_{\lambda}^{\prime}$ be a certain non-trivial character of $\mathrm{U}^{\prime}(F)$ that depends on $\psi$ and $\lambda$ (see Section 2.2 for its definition). In this paper, we prove the following local converse theorem for quasi-split metaplectic groups under the following hypothesis in $\operatorname{char}(F)=p\neq 2$ case. (There is no hypothesis in $\operatorname{char}(F)=0$ case.)
Working Hypothesis.
The $\gamma$-factors for $\widetilde{\mathrm{J}_{n}}\times\mathrm{GL}_{l}$ are properly defined in $\operatorname{char}(F)=p\neq 2$ cases. Furthermore, they satisfy natural properties of $\gamma$-factors. (For the precision of the natural properties, see (i)-(vii) except (vi) in Proposition 2.1 )
@@
Assume the above working hypothesis on $\operatorname{char}(F)=p\neq 2$ case. Let $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ be irreducible admissible $\mu_{\lambda}^{\prime}$-generic representations of $\widetilde{\mathrm{J}_{n}}(F)$, satisfying the condition
@@
Around 2017 Chai [Cha19] and Jacquet and Liu [JL18] independently refined Henniart’s result to its optimal form. Recently, this is also achieved by P. Yan, Q. Zhang [YZ23] in a different manner.
Recently, Q. Zhang [Q18, Q19] proved the supercuspidal cases of the conjecture for $\mathrm{Sp}_{2n}$ and $\mathrm{U}_{2n+1}$ using the theory of partial Bessel functions developed by Cogdell, Shahidi, and Tsai in [CST17]. In a similar vein, B. Liu and A. Hazeltine [LH22] established the conjecture for split even special orthogonal groups, while Y. Jo [Jo] refined Jiang and Soudry’s [JS03] result and extended Zhang’s [Q18] result to the generic cases. Notably, Jo’s work addressed local fields of positive characteristics as well as characteristic zero, in contrast to other LCT results that only dealt with characteristic zero cases. These proofs were purely local and did not invoke the weak Langlands functorial lifts to general linear groups, as in the recent results of [Cha19] and [JL18]. For other classical groups, Morimoto [M] demonstrated the theorem for even unitary groups by pulling back the recent results of [Cha19] and [JL18] using global (and hence) local descent methods.
@@
To successfully transfer the LCT, it is necessary to demonstrate that the LTC preserves certain properties and invariants of representations. Specifically, we consider genericity (Theorem 3.3) and $\gamma$-factors (Theorem 3.7). Indeed, these results can be found in [GS, Theorem 9.1] and [GS, Proposition 11.1, Corollary 11.2], respectively.
However, unfortunately, the proof of [GS, Theorem 9.1] is absent in the cited reference. To address this gap, we provide a complete proof demonstrating that the missing argument holds even in the case when $\operatorname{char}(F)=p$. Additionally, [GS, Proposition 11.1] relies on the general $\gamma$-factor theory, which is currently unavailable. Consequently, we furnish a comprehensive proof for this proposition using only the established theory on generic $\gamma$-factor (see Remark 3.8).
Theorem 3.3 and Theorem 3.7 form the main part of the paper and are discussed in Section 3, along with Section 4, which computes the twisted Jacquet module of the Weil representation—a crucial step in proving the preservation of genericity.
@@
When the residual characteristic of the $p$-adic field $F$ is odd, Zhang ([Q17]) directly proved this result under the condition that $\widetilde{\pi}$ and $\widetilde{\pi}^{\prime}$ share the same central characters. Our result covers cases where the residual characteristic of $F$ is 2 without requiring the assumption of identical central characters. Alternatively, we can remove the assumption on the same central characters from the local functional equation of doubling zeta integrals of metaplectic groups (see [GS, Section 4.3]).
@@ 2.3. $\gamma$-factors
Let $\mathrm{G}$ be either $\mathrm{H}_{m}=\mathrm{SO}_{2m+1}$ or $\widetilde{\mathrm{J}_{n}}=\widetilde{\mathrm{Sp}}_{2n}$ Let $\pi$ be an irreducible generic representation of $\mathrm{G}(F)$ and $\sigma$ an irreducible generic representation of $\mathrm{GL}_{l}(F)$. The local $\gamma$-factors can be defined in two distinct ways. The first approach is through the Langlands–Shahidi method ([Sha90], [L15], [Sz10]), while the second approach relies on the proportionality between the functional equations of local Rankin–Selberg type integrals (or Shimura type integrals) for $\mathrm{H}_{m}\times\mathrm{GL}_{l}$ (or $\widetilde{\mathrm{J}_{n}}\times\mathrm{GL}_{l}$). It is a product of a monomial of $t=q^{-s}$ and a rational function $\frac{Q(t)}{P(t)}$, where $Q(t),P(t)\in F[t]$ such that $Q(0)=P(0)=1$. For a precise definition, please refer to [So93] and [Kap15]. It is worth noting that the two definitions of $\gamma$-factors are reconciled by several properties, as explained in [Kap15]. Since we consider both the cases of $\operatorname{char}(F)=0$ and $\operatorname{char}(F)=p$ simultaneously, we adopt the definition of $\gamma$-factors from the Langlands–Shahidi method, as they are defined in both cases. It is worthy to mention that Langlands-Shahidi method produces not only $\gamma$-factors but also local $L$-factors as follows:
Write $\gamma(s,\pi\times\sigma,\psi)=\alpha\cdot t^{k}\cdot\frac{Q(t)}{P(t)}$ for some $\alpha\in F^{\times}$ and polynomials $Q(t),P(t)\in F[t]$ such that $\text{gcd}\big{(}Q(t),P(t)\big{)}=1$ and $Q(0)=P(0)=1$. Then $L_{\psi}(s,\pi\times\sigma,\psi)$ is defined by
(2.1) $L_{\psi}(s,\pi\times\sigma)\coloneqq Q(t)^{-1}$
We would like to recall some important properties of the $\gamma$-factors that will be utilized later in our discussion. (Caution : the property (vi) does not exist in the $\operatorname{char}(F)=p$ case.)
Proposition 2.1.
@@ 7. $(\mathrm{vii})$
    (Tempered $L$-function) Let $\pi,\sigma$ be irreducible tempered representations of $\mathrm{G}(F)$ and $\mathrm{GL}_{k}(F)$. Then $L_{\psi}(s,\pi\times\rho)$ is holomorphic for $\operatorname{Re}(s)>0$.
When $\mathrm{G}=\mathrm{H}_{m}\text{ and }\widetilde{\mathrm{J}_{n}}$, properties (i)-(vi) are proved in [Kap15, Theorem 1] (see also [Sha90, Theorem 3.5]) for $\operatorname{char}(F)=0$ case. However, property (vii) is only proved for $\mathrm{G}=\mathrm{H}_{m}$ ([CS, page 573], [HO13, Theorem 1.1]) in $\operatorname{char}{F}=0$ case and in [L15, Theorem 1.4] for $\operatorname{char}(F)=p$ case.
The property (vii) is proved in [CS, page 573], [HO13, Theorem 1.1] for $\operatorname{char}{F}=0$ case and in [CGL24, Theorem 1.1] for $\operatorname{char}{F}=p$ case.
@@
While the definition and standard properties of $\gamma$-factors for metaplectic groups in characteristic $\operatorname{char}(F)=p$ have not yet been firmly established, it is very reasonable to expect that the methods introduced in [L15] naturally extend to this case. However, due to the focus and limitations of this paper, we will adopt the working hypothesis that such extensions exist without delving into the details.
@@
The $\gamma$-factors for $\widetilde{\mathrm{J}_{n}}\times\mathrm{GL}_{l}$ are properly defined in $\operatorname{char}(F)=p$ cases. Furthermore, they satisfy properties (i)-(v) and (vii) in Proposition 2.1.
@@ 3. Local theta correspondence
For the rest of this paper, we focus exclusively on the case where $m=n$ and conveniently denote $V_{n}$ and $W_{n}$ as $V$ and $W$ respectively. Given $\widetilde{\pi}\in\operatorname{Irr}(\widetilde{\mathrm{J}_{n}})$, there exists a unique $\epsilon\in\{\pm 1\}$ such that $\theta_{\psi,V^{\epsilon},W}(\widetilde{\pi})$ is nonzero (see [GS, Theorem 1.4]).
@@ 3. $(\mathrm{iii})$
When $r=1$, the theorem is incorporated into [GI1, Theorem 11.5] in $\operatorname{char}(F)=0$ case. In the case of $\operatorname{char}(F)=p\neq 2$, a parallel reasoning to the proof of [GI1, Theorem 11.5] can be applied, utilizing the global-to-local argument from [GI1, subsection 11.6]. Notably, in this scenario, [GL, Theorem 1.1] plays a comparable role. Consequently, we may presume the validity of [GI1, Theorem 11.5] across all characteristics except 2.
@@
Let $\Xi$ be the global theta lift of $\Pi$ to $\mathrm{SO}(\mathbb{V})(\mathbb{A})$. Using [Fu, Proposition 3], we know that $\Xi$ is nonzero and generic. We claim that $\Xi$ is cuspidal. Suppose not, which means that there exists an integer $m<n$ and a $(2m+1)$-dimensional symmetric space $\mathbb{V}_{m}$ over $K$ such that the global theta lift $\Theta_{m}(\Pi)$ of $\Pi$ to $\mathrm{SO}(\mathbb{V}_{m})(\mathbb{A})$ is cuspidal. Let $\Theta_{m}(\Pi)_{0}$ be an irreducible summand of $\Theta_{m}(\Pi)$. Then $\Theta_{m}(\Pi)_{0,v_{0}}$ is the nonzero local theta lift of $\widetilde{\pi}$ to $\mathrm{SO}(V_{m})$. However, this contradicts the tower property of local theta lifts (see [Ku, Theorem 6.1]) since $\widetilde{\pi}$ is a supercuspidal representation of $\mathrm{H}_{n}(F)$. Therefore, we have shown that $\Xi$ is a cuspidal representation of $\mathrm{SO}(\mathbb{V})(\mathbb{A})$ with $\Xi_{v_{0}}=\theta_{\psi,V,W}(\widetilde{\pi})$.
@@ 5. The proof
For an irreducible generic tempered representation $\tau$ of $\mathrm{H}_{n}(F)$ and unitary supercuspidal representation $\sigma$ of $\mathrm{GL}_{k}(F)$, the followings hold.
@@ 1. $(\mathrm{i})$
    $\gamma(s,\tau\times\sigma,\psi)$ has no zero for $\operatorname{Re}(s)>0$.
@@ 2. $(\mathrm{ii})$
    If $\gamma(s,\tau\times\sigma,\psi)$ has a real pole at $s=s_{0}$, then the pole must be a simple pole at $s_{0}=1$ and $\sigma\simeq\sigma^{\vee}$.
@@
From the definition of local $L$-function (2.1), the zero of $\gamma(s,\tau\times\sigma,\psi)$ comes from the poles of $L(s,\tau\times\sigma)$. However, Proposition 2.1 (vii) implies that $L(s,\tau\times\sigma)$ has no pole for $\operatorname{Re}(s)>0$. Therefore $(\mathrm{i})$ of the theorem follows.
In the proof of [Jo, Proposition 4.12 $(\mathrm{iii})$], the requirement that $\pi$ is supercuspidal was imposed in order to apply the ‘Casselman–Shahidi’ Lemma ([Jo, Lemma 4.10]). However, it has been established in [Kim, Proposition 12.3] that the ‘Casselman–Shahidi Lemma’ also holds for tempered representations. Therefore, except for this specific assumption, all other arguments presented in [Jo, Proposition 4.12 $(\mathrm{iii})$] can be equally applied to tempered representations. Hence, property $(\mathrm{ii})$ is proven. ∎